OpenStudy (anonymous):

What are the domain of these two functions:

OpenStudy (anonymous):

\[y=-3\sec(\pi-2x)+5\]

OpenStudy (anonymous):

\[y=-2\sin(5/4x)\]

OpenStudy (anonymous):

@Nnesha

OpenStudy (anonymous):

somebody please help

OpenStudy (anonymous):

@Vocaloid

OpenStudy (anonymous):

Domain represents the value for x thus in the second equation is there any value that x cannot be?

OpenStudy (anonymous):

the sine curve has its domain unrestricted where as the secant function has vertical asymptotes, hence a restricted domain.

OpenStudy (anonymous):

therefore the domain for the first function is (-infinity,infinity)

OpenStudy (anonymous):

but I do not know the domain for the second function, what is it? @Deeezzzz

OpenStudy (anonymous):

Are you familiar with finding the period of the secant func?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

The period of the second function is 2

OpenStudy (anonymous):

@Deeezzzz

OpenStudy (mathstudent55):

The first function uses the secant function. Remember this identity. \(\sec \theta = \dfrac{1}{\cos \theta}\) The secant is not defined where the cosine equals zero.

OpenStudy (mathstudent55):

In the second function, you have the sine function. The sine function is defined for every value of theta.

OpenStudy (anonymous):

wait, so for the second function the domain is (-infinity,infinity)

OpenStudy (mathstudent55):

yes

OpenStudy (anonymous):

I still dont get the domain for the first function

OpenStudy (mathstudent55):

There is no restriction on the domain of the sine function.

OpenStudy (anonymous):

how would i write it down in terms of interval notation

OpenStudy (mathstudent55):

Where is the cosine equal to zero?

OpenStudy (anonymous):

its undefined?

OpenStudy (anonymous):

@mathstudent55

OpenStudy (mathstudent55):

The figure below shows where the cosine equals zero. It is at those points where the secant is undefined.